OpenStudy (anonymous):

Algebra 2: posting ....

OpenStudy (anonymous):

\[2(3n-4)/3=\frac{ 2 }{ 3 }(6n+5)\]

OpenStudy (anonymous):

distribute the 2 to the parenthesis\[{6n-8\over3}~~~={12n+10\over3}\]multiply both sides by 3 and the 3's cancel out\[6n-8=12n+10\]subtract 12n from both sides\[-6n-8=10\]add 8 to both sides\[-6n=18\]divide both sides by -6\[n=~?\]

OpenStudy (anonymous):

@NeedMathHelpBad what did you get for 'n'? :)

OpenStudy (anonymous):

-3?

OpenStudy (anonymous):

positutely! :DD

OpenStudy (anonymous):

thank ya

OpenStudy (anonymous):

anytime! :)